solve and check for extraneous solutions: natural logarithms (In)square root (x-3)=1
Answer by edjones(8007)

In(sqrt(x-3))=1
e^(In(sqrt(x-3)))=e^1
sqrt(x-3)=e
x-3=e^2 square each side.
x=e^2+3
.
Ed